# Learning Goal-Conditioned Policies from Sub-Optimal Offline Data via Metric Learning

### v2

@@ 3.4 Practical Implementation
Stabilizing the loss In practice, we take the logarithm of the contrastive term (second term) in Equation 1. We have found this formulation to stabilize the training, in particular for environments with longer episode horizons:
$\mathcal{L}_{\theta}(D)=\operatorname{\mathbb{E}}_{D}\left[(\left\lVert\phi_{\theta}(s^{\prime})-\phi_{\theta}(s)\right\rVert_{2}-1)^{2}-\lambda\log\left\lVert\phi_{\theta}(s^{\prime\prime})-\phi_{\theta}(s)\right\rVert_{2}\right].$ (4)
Learning with images When learning with images, often the goal information is present in the image, which can break the connectivity assumption: two images with different goal positions are not connected by any path. However, in the case of finite MDPs, we can easily recover it by introducing an additional super-state that connects every termination state together. We present the implementation details of such a super-state in Appendix A.5. The introduction of this super-state and the consequent connection of the environments with different goals leads to a modification of the learned representation. Figure 5 shows this learned representation. On the left is the distribution of the states for one single goal, while on the right is the distribution of all the goals connected by the super-state (red star). The solution to this representation is a radial distribution of the states connected in the middle by the super-state. All the states with the same goal (orange dots in the image) compose one ray of the overall distribution.
@@
4 Benchmark Evaluation
We evaluate MetricRL against standard baselines in offline reinforcement learning across multiple environments. In all experiments we consider three types of offline datasets: a low dataset, often collected using a random policy or an untrained agent; a medium dataset, collected using the policy of an online RL agent during training or adding stochasticity to a fully trained agent, and a high dataset, collected using the policy of a fully-trained online RL agent.
For baselines, we consider:
  • CQL [26] introduces a conservative term in the estimation of the Q value. This term penalizes the highest values of the estimated Q function;
  • BCQ [27] perturbs the policy learned with a VAE with a DDPG term;
  • BEAR [28] constrains the learned policy to the behavioral one estimated with BC;
  • PLAS [29] trains the policy within the latent space of a conditional VAE trained on the Offline dataset;
  • IQL [15] avoids sampling out-of-distribution actions using a SARSA like critic update with quantile regression;
  • ContrastiveRL [12] approximates the value function of the policy that collected the dataset using contrastive learning. To adapt it to offline RL, the objective is coupled with a behavioral cloning term.
  • QRL [13] estimates a quasimetric to approximate the value function using a contrastive learning formulation. The model is paired with a learned policy regularized with a behavioral cloning term to avoid out-of-distribution state-actions pairs in offline RL conditions.
  • GoFAR [14] estimates an offline goal-conditioned RL policy by recasting it as a state occupancy matching problem.
For each model, we perform standard hyperparameter tuning or use the author’s suggested hyperparameters (if available). The complete list of training hyperparameters is available in Appendix A.6, CQL, BCQ, BEAR, PLAS, IQL are implemented using Seno and Imai [30]. For the remaining models we use the author’s provided code.
@@ 4 Benchmark Evaluation
  • Maze2D [31]: a navigation task within a two-dimensional maze, with continuous actions and Newtonian physics. We consider three sizes of the maze: u-maze, medium and large. For each size, we use a uniform random policy to collect the low dataset, a policy with Ornstein-Uhlenbeck noise [32] to collect the medium dataset, and we use the Minari dataset provided in D4RL for the high dataset [31];
  • Reach [33]: a manipulation task with a 7-DoF robot with continuous actions to reach a randomly-selected goal position in the workspace. To collect the datasets we employ a PPO agent trained on dense rewards along three different stages of training. For the low dataset, we use the policy of the randomly-initialized agent, for the medium dataset we use a PPO agent achieving half of the optimal expected reward, and for the high dataset we use the policy of the fully-converged agent;
@@
  • Minigrid [1]: a navigation task on a grid-like 2D room with discrete actions. We restrict the action space of the agent to navigation actions and remove rotations. To collect the offline datasets we employ the same strategy as above, that is $\epsilon$-greedy on a fully trained DQN agent. We consider 2 tasks, Minigrid Empty consists of an open grid with external walls as the only obstacles. Minigrid DoorKey is a three-step task where the agent must first pick-up a key, then open a door, then reach a goal position;
@@
MetricRL provides a stabler training across different data distributions in offline datasets: We conduct an additional experiment on the Minigrid DoorKey environment and consider a more fine discretization of the distribution of rewards. We collect multiple datasets using an increasing value of $\epsilon$ for an $\epsilon$-greedy optimal DQN agent. As shown in Figure 1, MetricRL (red line) requires datasets with a smaller average reward than the baselines to achieve optimal performance. Additionally, the results show that MetricRL does not suffer from out-of-distribution data when the dataset has a very narrow distribution (almost all or all optimal trajectories).
@@
MetricRL can learn from sub-optimal datasets of images: To evaluate the performance of MetricRL when provided with high-dimensional observations (images) we reuse the MiniGrid Empty and DoorKey environments and introduce a super-state following the discussion in Section 3.4. For every model, we introduce a CNN architecture that maps the images to a lower-dimensional representation. This highlights that MetricRL maintains its performance, and remains the only model able to learn optimal behavior from sub-optimal datasets. We present an additional result in the Appendix with Figure 14.
4.1 Additional Results
In Appendix A.7 we explore the sample efficiency of MetricRL in scenarios with large state spaces. The results show that our method significantly outperforms temporal-difference (TD) methods (e.g., DQN [34]): to solve larger state space problems, we require a linear increase in the number of training iterations against the exponential increase of TD methods. In Appendix A.8, we show how MetricRL can be used in multi-goal tasks without modifications, considering changing multiple goals and discount factors at execution time.
5 Discussion
MetricRL recasts the computation of the value function as a problem of measuring distances in an appropriate learned metric space. To do so, it requires two additional assumptions on the MDPs it is applied to: the existence of inverse actions and the connectivity of the dataset used to learn the value function. As stated in Section 3.4, the connectivity assumption can be solved using “super-states” to join the states into a unique connected component. The existence of inverse actions, on the other hand, defines a trade-off. It limits the applicability of the proposed method to MDPs where the assumption is respected. On the other hand, it greatly simplifies the learning process by imposing a relevant bias on the representation. MetricRL, in fact, requires significantly fewer data points as it can interpolate missing transitions when the inverse transition is present in the data. This allows to estimate effective policies even in severely sub-optimal data conditions as shown in Section 4.
@@ 6 Related Work
Offline RL: The challenge of learning a policy from a static sub-optimal dataset of transitions and rewards has been extensively studied [7]. The problem of exploration is not considered as it is assumed that the dataset given contains all the relevant information to estimate an optimal policy. Methods can be roughly divided into four main categories. The first one is constraining the learned policy to not deviate much from the policy that collected the dataset: Kumar et al. [28] restrict policies to have the same support as the behavior policy rather than the policy itself; Zhou et al. [29] implicitly constrain the policy by learning it using latent representations of actions with Gaussian prior (VAE), the constraint given by the KL divergence term; Siegel et al. [35] learns the behavioral policy explicitly. The second category introduces a penalty in the reward function based on the uncertainty of the transitions or the reward function. This penalty has been defined as the uncertainty of the learned $Q$ function, [36, 37], or a measure of pessimism (regularization of the highest $Q$) [26]. A third family of methods instead uses model-based RL and explicitly computes a model of the environment that can be used in different forms to regularize the learned policy [38, 39, 40, 27]. The last category includes in-sample algorithms which restrict the learning of the policy only on data within the provided dataset [15]. Our method falls in this last category, as can be seen in Equation 3. A major challenge in all of these methods is their effectiveness when the quality of the data decreases substantially. In this work we propose a method that can estimate a high-return policy independently from the distribution of the data used.
@@
Contrastive Learning: Representation learning techniques have been used to aid the RL problem. Several works have used contrastive learning models to speed up or improve the generalization of a classic RL algorithm, [41, 42, 43, 44]. Other applications include reward function estimation from demonstrations [45, 46] or generating intermediate goals for curriculum learning [47]. In [12, 48] contrastive learning is used to estimate the discounted state occupancy measure which is equivalent to the Q function in some particular cases. This method however works only when the reward function can be expressed as a goal reaching density and doesn’t estimate the optimal Q value but rather the Q value of a current policy, thus still requiring the concurrent learning of a policy in a classic RL fashion. Zhu et al. [49] proposes the use of contrastive learning to build a representation of states where distances are correlated with reachability in terms of actions. After the representation is learned, they propose to explicitly build the graph of the offline dataset and apply value iteration to get an estimate of the value function. The policy can be obtained by applying Dijkstra on the learned graph. More similar to this work [13, 19] estimate the optimal value function rather than the policy one. In Yang et al. [19] the authors use contrastive learning to find a representation where connections between states in action terms can be estimated in terms of Euclidean distances. The value function can then be recovered as the sum of the shortest path distances between the goal and the current state using depth-first search classical algorithms. Wang et al. [13] learns a map between pairs of states and a quasimetric representing the estimated optimal value function of a goal-conditioned MDP. This is done by setting the distance between two consecutive states to be equal to the reward between them and the distance between random pairs of states to be maximized. The optimal value function can then be defined as the distance between each state and the goal state. While being more general, quasimetrics cannot capture the appropriate bias induced by the inverse actions assumption. We show empirically that our proposed method is more effective in estimating a policy when data is severely sub-optimal. In this paper, we describe a property of representations needed to ensure the optimality of the critic function and propose a simple method to recover such a representation for a particular class of MDPs. Moreover, differently from the works above, we study the effectiveness of this methodology in handling offline datasets collected by policies of different values.
@@ 7 Conclusions
In this paper, we proposed MetricRL, a novel approximation method for the optimal value function of sparse, deterministic, goal-conditioned MDPs. MetricRL relies on learning a distance monotonic representation of the state space, allowing it to define a value function that is correlated with the distance of each state to the goal. We have proved that, when the representation is indeed distance monotonic, a greedy policy on this approximated value function is optimal for the class of MDPs stated above. Experimentally, we have shown that MetricRL outperforms prior offline RL methods in learning near-optimal behavior from severely sub-optimal datasets. For future work, we plan on generalizing the notion of distance monotonicity to quasimetrics [50, 51], extend our method to stochastic MDPs and adapt it to online reinforcement learning problems.
8 Acknowledgements
@@ Appendix A Appendix
A.1 Additional Related Work
Other metrics: Other measures of state similarity from a control perspective have been explored before. Bisimulation metric defines a measure of similarity between states in terms of future transitions and rewards, [52, 53]. These methods are theoretically grounded but particularly difficult to make them work in practice. This is especially true in the case of continuous spaces. Older work has explored different forms of value function approximation. By parametrically approximating the map between each state and its value, [54] approximates a notion of similarity (in a value sense) between states with an appropriate kernel and rewrite the Bellman operator as a function of this kernel. This still needs to solve the optimization problem with value iteration techniques. Proto-value functions, [55], instead express the transition and reward functions as linear matrices, the value function problem has exact solutions and can be estimated with an appropriate kernel method at the cost of expressivity.
A.2 Sparse Goal-Conditioned Value Functions in Isometric Spaces
Recall the definition of the optimal value function described above as the maximum over the possible policies of the expectation of the discounted cumulative reward: ${V^{*}(s,s_{g})=\max_{\pi}\operatorname{\mathbb{E}}_{\pi}\left[\sum_{t}\gamma^{t}r_{t}|s_{0}=s,a=\pi(s,s_{g})\right]}$.
@@ A.2 Sparse Goal-Conditioned Value Functions in Isometric Spaces
Assuming $\phi$ is an isometry, ${d_{G}(s,s_{g})=d_{Z}(\phi(s),\phi(s^{\prime}))}$ resulting in the equivalency: ${V^{*}(s,s_{g})=\gamma^{d_{Z}(\phi(s),\phi(s^{\prime}))}r_{g}}$.
A.3 Distance Monotonicity Measure

### v3

@@ 3.4 Practical Implementation
Stabilizing the loss In practice, the negative term of the loss 1 is unbounded. In practice, we propose to take the logarithm of this contrastive term. This effectively changes the dynamics of the learning, resulting in a weaker pull from the negative terms. We have found this formulation to stabilize the training, in particular for environments with longer episode horizons:
$\mathcal{L}_{\theta}(D)=\mathbb{E}_{D}\left[(\left\lVert\phi_{\theta}(s^{\prime})-\phi_{\theta}(s)\right\rVert_{2}-1)^{2}-\lambda\log\left\lVert\phi_{\theta}(s^{\prime\prime})-\phi_{\theta}(s)\right\rVert_{2}\right].$ (4)
Learning with images When learning with images, often the goal information is present in the image, e.g., the position of the green square in the grid experiment in Figure 5, which can break the connectivity assumption: two images with different goal positions are not connected by any path. This can be a problem in practice as the second term in Equation 1 can grow indefinitely when comparing representations of images with different goals. However, in the case of finite MDPs, we can easily recover it by introducing an additional super-state that connects every termination state together. We present the implementation details of such a super-state in Appendix A.4. The introduction of this super-state and the consequent connection of the environments with different goals leads to a modification of the learned representation. Figure 5 shows this learned representation. On the left is the distribution of the states for one single goal, while on the right is the distribution of all the goals connected by the super-state (red star). The solution to this representation is a radial distribution of the states connected in the middle by the super-state. All the states with the same goal (orange dots in the image) compose one ray of the overall distribution.
@@
4 Results
We evaluate MetricRL against state-of-the-art baselines in offline reinforcement learning across multiple environments. In all experiments we consider three types of offline datasets: a low dataset, often collected using a random policy or an untrained agent; a medium dataset, collected using the policy of an online RL agent during training or adding stochasticity to a fully trained agent, and a high dataset, collected using the policy of a fully-trained online RL agent.
For baselines, we consider the following: CQL (Kumar et al., 2020), BCQ (Fujimoto et al., 2019), BEAR (Kumar et al., 2019), PLAS (Zhou et al., 2021), IQL (Kostrikov et al., 2021), ContrastiveRL (Eysenbach et al., 2022), QRL (Wang et al., 2023), GoFAR (Ma et al., 2022b), HIQL (Park et al., 2023a). More details in Appendix A.6.
For each model, we perform standard hyperparameter tuning or use the author’s suggested hyperparameters (if available). The complete list of training hyperparameters is available in Appendix A.6, CQL, BCQ, BEAR, PLAS, IQL are implemented using Seno and Imai (2022). For the remaining models, we use the author’s provided code.
@@ 4 Results
  • Maze2D (Fu et al., 2020): a navigation task within a two-dimensional maze, with continuous actions and Newtonian physics. We consider three sizes of the maze: u-maze, medium and large. For each size, we use a uniform random policy to collect the low dataset, a policy with Ornstein-Uhlenbeck noise (Uhlenbeck and Ornstein, 1930) to collect the medium dataset, and we use the Minari dataset provided in D4RL for the high dataset (Fu et al., 2020);
  • Reach (Plappert et al., 2018): a manipulation task with a 7-DoF robot with continuous actions to reach a randomly-selected goal position in the workspace. To collect the datasets we employ a PPO agent trained on dense rewards along three different stages of training. For the low dataset, we use the policy of the randomly-initialized agent, for the medium dataset we use a PPO agent achieving half of the optimal expected reward, and for the high dataset we use the policy of the fully-converged agent;
@@
  • Minigrid (Chevalier-Boisvert et al., 2023): a navigation task on a grid-like 2D room with discrete actions. We restrict the action space of the agent to navigation actions and remove rotations. To collect the offline datasets we employ the same strategy as above, that is $\epsilon$-greedy on a fully trained DQN agent. We consider 2 tasks, Minigrid Empty consists of an open grid with external walls as the only obstacles. Minigrid DoorKey is a three-step task where the agent must first pick-up a key, then open a door, then reach a goal position;
@@ 4.1 Discussion
MetricRL provides more stable training across different data distributions in offline datasets: We conduct an additional experiment on the Minigrid DoorKey environment and consider a finer discretization of the distribution of rewards. We collect multiple datasets using an increasing value of $\epsilon$ for an $\epsilon$-greedy optimal DQN agent. As shown in Figure 1, MetricRL (red line) requires datasets with a smaller average reward than the baselines to achieve optimal performance. Additionally, the results show that MetricRL does not suffer from out-of-distribution data when the dataset has a very narrow distribution (almost all or all optimal trajectories).
@@
MetricRL can learn from sub-optimal datasets of images: To evaluate the performance of MetricRL when provided with high-dimensional observations (images) we reuse the MiniGrid Empty and DoorKey environments and introduce a super-state following the discussion in Section 3.4. For every model, we introduce a CNN architecture that maps the images to a lower-dimensional representation. This highlights that MetricRL maintains its performance, and remains the only model able to learn optimal behavior from sub-optimal datasets. We present additional results in line with this on multiple environments, see Figure 15 in Appendix.
In Appendix A.7 we explore the sample efficiency of MetricRL in scenarios with large state spaces. The results show that our method significantly outperforms temporal-difference (TD) methods (e.g., DQN (Mnih et al., 2015)): to solve larger state space problems, we require a linear increase in the number of training iterations against the exponential increase of TD methods. In Appendix A.8, we show how MetricRL can be used in multi-goal tasks without modifications, considering changing multiple goals and discount factors at execution time.
4.2 Metric Space Ablation
Table 1: Ablation on symmetrizing data in the Maze2D Large environment.
Dataset QRL QRL - sym MetricRL
Low 0.13 $\pm$ 0.05 0.25 $\pm$ 0.04 0.51 $\pm$ 0.05
Medium 0.22 $\pm$ 0.03 0.52 $\pm$ 0.02 0.56 $\pm$ 0.05
High 0.28 $\pm$ 0.08 0.56 $\pm$ 0.12 0.60 $\pm$ 0.09
We further ablate the impact of the symmetry bias and the use of a Euclidean embedding for value function approximation. Specifically, we compare MetricRL against an equivalent model that employs a quasimetric critic, following Wang et al. (2023). For a fair comparison, we use the same actor optimization as in Equation 3. We also evaluate a symmetrized variant of the quasimetric model by augmenting the dataset with synthetic reverse transitions, i.e., $(s^{\prime},a,s)$. Results on the Maze2D-large environment (Table 1) show that incorporating symmetry can significantly enhance performance when it reflects a valid inductive bias. Moreover, Euclidean embeddings often prove easier to optimize, particularly in low-quality or sub-optimal data regimes.
4.3 Limitations
MetricRL recasts the computation of the value function as a problem of measuring distances in an appropriate learned metric space. To do so, it requires two additional assumptions on the MDPs it is applied to: the existence of inverse actions and the connectivity of the dataset used to learn the value function. As stated in Section 3.4, the connectivity assumption can be solved using “super-states” to join the states into a unique connected component. The existence of inverse actions, on the other hand, defines a trade-off. It limits the applicability of the proposed method to MDPs where the assumption is respected. On the other hand, it greatly simplifies the learning process by imposing a relevant bias on the representation. MetricRL, in fact, requires significantly fewer data points as it can interpolate missing transitions when the inverse transition is present in the data. This allows to estimate effective policies even in severely sub-optimal data conditions.
@@ 5 Related Work
Offline RL: The challenge of learning a policy from a static sub-optimal dataset of transitions and rewards has been extensively studied (Levine et al., 2020). The problem of exploration is not considered as it is assumed that the dataset given contains all the relevant information to estimate an optimal policy. Methods can be roughly divided into four main categories. The first one is constraining the learned policy to not deviate much from the policy that collected the dataset: Kumar et al. (2019) restrict policies to have the same support as the behavior policy rather than the policy itself; Zhou et al. (2021) implicitly constrain the policy by learning it using latent representations of actions with Gaussian prior (VAE), the constraint given by the KL divergence term; Siegel et al. (2020) learns the behavioral policy explicitly. The second category introduces a penalty in the reward function based on the uncertainty of the transitions or the reward function. This penalty has been defined as the uncertainty of the learned $Q$ function, (Kidambi et al., 2020; Yu et al., 2020), or a measure of pessimism (regularization of the highest QQ(Kumar et al., 2020). A third family of methods instead uses model-based RL and explicitly computes a model of the environment that can be used in different forms to regularize the learned policy (Matsushima et al., 2020; Yu et al., 2021; Rigter et al., 2022; Fujimoto et al., 2019). The last category includes in-sample algorithms which restrict the learning of the policy only on data within the provided dataset and reweighted by an estimate of the advantage function. This family of methods has been referred to as weighted supervised learning (Wang et al., 2018) or maximum likelihood (Nair et al., 2020). It has been extended with different forms of regularization including expectile regression for the critic (Kostrikov et al., 2021), penalizing out-of-distribution actions in the critic (Chebotar et al., 2021), trust regions (Mao et al., 2023), goal relabeling (Yang et al., 2022) and Generative Adversarial Networks (Wang et al., 2024). The policy estimate of our proposed method falls in this last category, as can be seen in Equation 3. The use of Temporal Difference learning for the critic, however, subjects these methods to the problem of distribution shift in Offline RL. In this work, we propose a method that can estimate a high-return policy independently from the distribution of the data used even when the quality of the data decreases substantially.
@@
Contrastive Learning: Representation learning techniques have been used to aid the RL problem. Several works have used contrastive learning models to speed up or improve the generalization of a classic RL algorithm, (Laskin et al., 2020; Oord et al., 2018; Anand et al., 2019; Stooke et al., 2021). Other applications include reward function estimation from demonstrations (Ma et al., 2022a; Sermanet et al., 2018) or generating intermediate goals for curriculum learning (Venkattaramanujam et al., 2019). In Eysenbach et al. (2022); Hatch et al. (2023) contrastive learning is used to estimate the discounted state occupancy measure which is equivalent to the Q function in some particular cases. This method however works only when the reward function can be expressed as a goal reaching density and doesn’t estimate the optimal Q value but rather the Q value of a current policy, thus still requiring the concurrent learning of a policy in a classic RL fashion. Zhu et al. (2022) proposes the use of contrastive learning to build a representation of states where distances are correlated with reachability in terms of actions. After the representation is learned, they propose to explicitly build the graph of the offline dataset and apply value iteration to get an estimate of the value function. The policy can be obtained by applying Dijkstra on the learned graph. More similar to this work Wang et al. (2023); Yang et al. (2020) estimate the optimal value function rather than the policy one. In Yang et al. (2020) the authors use contrastive learning to find a representation where connections between states in action terms can be estimated in terms of Euclidean distances. The value function can then be recovered as the sum of the shortest path distances between the goal and the current state using classical depth-first search algorithms. Wang et al. (2023) learns a map between pairs of states and a quasimetric representing the estimated optimal value function of a goal-conditioned MDP. This is done by setting the distance between two consecutive states to be equal to the reward between them and the distance between random pairs of states to be maximized. The optimal value function can then be defined as the distance between each state and the goal state. While being more general, quasimetrics cannot capture the appropriate bias induced by the inverse actions assumption. We show empirically that our proposed method is more effective in estimating a policy when data is severely sub-optimal. In this paper, we describe a property of representations needed to ensure the optimality of the critic function and propose a simple method to recover such a representation for a particular class of MDPs. Moreover, differently from the works above, we study the effectiveness of this methodology in handling offline datasets collected by policies that are not necessarily optimal.
Other metrics: Other measures of state similarity from a control perspective have been explored before. Bisimulation metric defines a measure of similarity between states in terms of future transitions and rewards, (Ferns and Precup, 2014; Castro, 2020). These methods are theoretically grounded but particularly difficult to make them work in practice. This is especially true in the case of continuous spaces. Older work has explored different forms of value function approximation. By parametrically approximating the map between each state and its value, (Ormoneit and Sen, 2002) approximates a notion of similarity (in a value sense) between states with an appropriate kernel and rewrite the Bellman operator as a function of this kernel. This still needs to solve the optimization problem with value iteration techniques. Proto-value functions, (Parr et al., 2008), instead express the transition and reward functions as linear matrices, the value function problem has exact solutions and can be estimated with an appropriate kernel method at the cost of expressivity.
@@ 6 Conclusions
In this paper, we proposed MetricRL, a novel approximation method for the optimal value function of sparse, deterministic, goal-conditioned MDPs. MetricRL relies on learning a distance monotonic representation of the state space, allowing it to define a value function that is correlated with the distance of each state to the goal. We have proved that, when the representation is indeed distance monotonic, a greedy policy on this approximated value function is optimal for the class of MDPs stated above. Experimentally, we have shown that MetricRL outperforms prior offline RL methods in learning near-optimal behavior from severely sub-optimal datasets. For future work, we plan on generalizing the notion of distance monotonicity to quasimetrics (Durugkar et al., 2021; Durugkar, 2023), extending our method to stochastic MDPs and adapting it to online reinforcement learning problems.
Acknowledgments
@@ Appendix A Appendix
A.1 Sparse Goal-Conditioned Value Functions in Isometric Spaces
Recall the definition of the optimal value function described above as the maximum over the possible policies of the expectation of the discounted cumulative reward: $V^{*}(s,s_{g})=\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t}\gamma^{t}r_{t}|s_{0}=s,a=\pi(s,s_{g})\right]$}.
@@ A.1 Sparse Goal-Conditioned Value Functions in Isometric Spaces
Assuming $\phi$ is an isometry, $d_{G}(s,s_{g})=d_{Z}(\phi(s),\phi(s^{\prime}))$} resulting in: $V^{*}(s,s_{g})=\gamma^{d_{Z}(\phi(s),\phi(s^{\prime}))}r_{g}$}.
A.2 Distance Monotonicity Measure